# SIRD: A Sender-Informed, Receiver-Driven Datacenter Transport Protocol

### v4

@@ 8. Conclusion
SIRD is a novel congestion control protocol that tackles the fundamental tension of receiver-driven designs which is the management of shared links. SIRD applies proactive scheduling to exclusive receiver downlinks, which are statistically the most congested and reactively allocates the bandwidth of shared links. SIRD achieves high link utilization with hardly any buffering and is the only protocol, among the six evaluated, that robustly and simultaneously delivers high throughput with minimal queuing and low latency. SIRD does not depend on advanced switch ASIC features nor Ethernet priorities, which makes it pragmatically deployable on existing and heterogeneous datacenter networks.
@@ Appendix A Appendix
Table 3. ASIC bisection bandwidth (in Tbps) and buffer sizes (in MB) data for Figure 1. Note that buffer architectures differ regarding the extent to which they are shared between ports.
@@
Figure 11. Message latency slowdown for the median and the $99^{th}$ percentile at 50% offered application load. Each bar group contains the messages in the following size ranges: $0\leq A<MSS\leq B<1\times BDP\leq C<8\times BDP\leq D$.
@@
Normalized 99th percentile slowdown of all messages at 50% load.
DCTCP 0.9 0.98 0.98 0.9 0.94 0.94 0.81 0.83 0.96 0.92 0.17
Swift 0.89 0.92 0.91 0.87 0.92 0.89 0.92 1.0 0.92 0.92 0.13
ExpressPass 0.46 0.93 0.93 0.45 0.89 0.88 0.48 0.94 0.95 0.77 0.5
Homa 1.0 1.0 0.97 0.99 0.96 0.99 1.0 0.99 0.98 0.99 0.04
dcPIM 0.71 0.92 1.0 0.74 1.0 1.0 0.71 0.92 1.0 0.89 0.29
SIRD 0.96 0.97 0.98 1.0 0.98 0.96 0.96 0.96 1.0 0.97 0.04
Normalized maximum goodput across applied load levels.
DCTCP unstable 85.81 9.71 3.68 62.13 7.94 unstable unstable 260.51 71.63 256.83
Swift unstable unstable 8.41 unstable unstable 10.12 unstable unstable 125.77 48.1 117.36
ExpressPass 1.0 1.0 1.0 unstable 1.0 1.0 1.0 1.0 1.0 1.0 0.0
Homa 137.45 115.99 30.18 7.06 434.71 79.62 143.38 150.57 145.29 138.25 427.65
dcPIM 52.07 7.76 3.1 1.04 51.24 13.09 34.25 7.18 7.41 19.68 51.03
SIRD 12.05 9.91 2.68 1.0 38.9 7.53 9.37 10.03 10.27 11.3 37.9
Normalized maximum ToR queuing across applied load levels.
Table 4. Normalized data used in Figure 4. Performance is normalized to the best performing protocol on each experiment. Experiments in which the protocol is unable to deliver the specified throughput or network queuing grows infinitely (unstable) are excluded from the calculation of mean and range.
Config Default Core Incast mean range
Wload WKa WKb WKc WKa WKb WKc WKa WKb WKc
DCTCP 9.92 7.9 6.91 5.97 5.71 5.54 unstable 71.06 32.06 18.13 65.52
Swift 8.61 7.37 10.56 5.85 5.68 7.58 68.21 38.51 20.69 19.23 62.53
ExpressPass unstable 36.94 29.26 unstable 28.56 24.42 unstable 47.45 47.45 35.68 23.03
Homa 1.29 2.47 3.62 1.19 1.8 2.15 1.27 5.36 7.37 2.95 6.18
dcPIM 2.16 7.37 6.66 1.88 6.2 5.76 2.11 7.7 11.48 5.7 9.6
SIRD 1.7 3.53 3.99 1.37 2.28 2.54 1.65 3.76 5.79 2.96 4.42
99th percentile slowdown of all messages at 50% load.
DCTCP 74.65 83.85 83.95 43.63 48.54 47.42 67.8 70.53 80.74 66.79 40.32
Swift 74.52 78.16 78.69 42.16 47.41 45.15 76.8 85.46 77.57 67.32 43.3
ExpressPass 38.04 79.68 80.42 21.78 45.83 44.42 40.26 80.1 80.1 56.74 58.64
Homa 83.39 85.23 83.55 47.73 49.73 50.02 83.39 84.87 82.79 72.3 37.5
dcPIM 58.94 78.42 86.03 35.91 51.68 50.59 59.61 79.02 84.26 64.94 50.12
SIRD 79.74 82.27 84.71 48.27 50.47 48.75 79.7 81.98 84.13 71.11 36.44
Maximum goodput across applied load levels (Gbps).
DCTCP unstable 7.0 2.7 8.3 2.67 1.46 unstable unstable 21.06 7.2 19.6
Swift unstable unstable 2.33 unstable unstable 1.87 unstable unstable 10.17 4.79 8.3
ExpressPass 0.06 0.08 0.28 unstable 0.04 0.18 0.08 0.08 0.08 0.11 0.24
Homa 8.63 9.46 8.37 15.91 18.68 14.68 12.17 12.17 11.75 12.42 10.31
dcPIM 3.27 0.63 0.86 2.34 2.2 2.41 2.91 0.58 0.6 1.76 2.69
SIRD 0.76 0.81 0.75 2.26 1.67 1.39 0.79 0.81 0.83 1.12 1.51
Maximum ToR queuing across applied load levels (MB).
Table 5. Raw data used in Figure 4. Experiments in which the protocol is unable to deliver the specified throughput or network queuing grows infinitely (unstable) are excluded from the calculation of mean and range.
@@
Figure 12. Mean ToR queuing vs. achieved goodput. Configurations: Balanced (top), Core (middle), Incast (bottom).

### v5

@@ 8 Conclusion
SIRD takes a holistic approach to datacenter congestion management by tackling the fundamental tension of receiver-driven designs, which is the management of shared links. SIRD applies proactive scheduling to exclusive receiver downlinks, which are statistically the most congested, and leverages sender and fabric congestion feedback to reactively allocate the bandwidth of shared links. Compared to existing designs, SIRD manages to simultaneously deliver high link utilization with hardly any network buffering, and low message latency. SIRD does so without assuming advanced switch ASIC features nor switch priority queues, which makes it compatible with existing and heterogeneous datacenter networks. We implemented a prototype of SIRD in software, on the Caladan stack, and showed that it can efficiently allocate credit while delivering 100Gbps.
Acknowledgements
The authors thank John Ousterhout for numerous discussions on the topic of receiver-driven protocols. We thank our shepherd Srikanth Sundaresan, Nate Foster, Katerina Argyraki, Haitham Al Hassanieh, James Larus, Charly Castes, Neelu Kalani, Boris Pismenny, Rui Yang, Mahyar Emami, Sahand Kashani, Mia Primorac, Francois Costa, and all anonymous reviewers for their insightful comments.
This work was funded in part by the Microsoft-EPFL Joint Research Center and used CloudLab [27] for 100GbE experiments. Ryan Kosta participated in EPFL’s Excellence Research Internship Program.
@@ Appendix A Appendix
Table 3: ASIC bisection bandwidth (in Tbps) and buffer sizes (in MB). Note that buffer architectures differ regarding the extent to which they are shared between ports.
@@
Figure 12: Median and $99^{th}$ percentile slowdown at 50% offered application load. Each bar group contains the messages in the following size ranges: $0\leq A<MSS\leq B<1\times BDP\leq C<8\times BDP\leq D$.
@@
Normalized 99th percentile slowdown of all messages at 50% load.      DCTCP 0.9 0.98 0.98 0.9 0.94 0.94 0.81 0.83 0.96 0.92 0.17 Swift 0.89 0.92 0.91 0.87 0.92 0.89 0.92 1.0 0.92 0.92 0.13 ExpressPass 0.46 0.93 0.93 0.45 0.89 0.88 0.48 0.94 0.95 0.77 0.5 Homa 1.0 1.0 0.97 0.99 0.96 0.99 1.0 0.99 0.98 0.99 0.04 dcPIM 0.71 0.92 1.0 0.74 1.0 1.0 0.71 0.92 1.0 0.89 0.29 SIRD 0.96 0.97 0.98 1.0 0.98 0.96 0.96 0.96 1.0 0.97 0.04    Normalized maximum goodput across applied load levels.      DCTCP unstable 85.81 9.71 3.68 62.13 7.94 unstable unstable 260.51 71.63 256.83 Swift unstable unstable 8.41 unstable unstable 10.12 unstable unstable 125.77 48.1 117.36 ExpressPass 1.0 1.0 1.0 unstable 1.0 1.0 1.0 1.0 1.0 1.0 0.0 Homa 137.45 115.99 30.18 7.06 434.71 79.62 143.38 150.57 145.29 138.25 427.65 dcPIM 52.07 7.76 3.1 1.04 51.24 13.09 34.25 7.18 7.41 19.68 51.03 SIRD 12.05 9.91 2.68 1.0 38.9 7.53 9.37 10.03 10.27 11.3 37.9    Normalized maximum ToR queuing across applied load levels.
Table 4: Normalized data used in Figure 5. Performance is normalized to the best performing protocol on each experiment. Experiments in which the protocol is unable to deliver the specified throughput or network queuing grows infinitely (unstable) are excluded from the calculation of mean and range.
DCTCP 74.65 83.85 83.95 43.63 48.54 47.42 67.8 70.53 80.74 66.79 40.32 Swift 74.52 78.16 78.69 42.16 47.41 45.15 76.8 85.46 77.57 67.32 43.3 ExpressPass 38.04 79.68 80.42 21.78 45.83 44.42 40.26 80.1 80.1 56.74 58.64 Homa 83.39 85.23 83.55 47.73 49.73 50.02 83.39 84.87 82.79 72.3 37.5 dcPIM 58.94 78.42 86.03 35.91 51.68 50.59 59.61 79.02 84.26 64.94 50.12 SIRD 79.74 82.27 84.71 48.27 50.47 48.75 79.7 81.98 84.13 71.11 36.44    Maximum goodput across applied load levels (Gbps).      DCTCP unstable 7.0 2.7 8.3 2.67 1.46 unstable unstable 21.06 7.2 19.6 Swift unstable unstable 2.33 unstable unstable 1.87 unstable unstable 10.17 4.79 8.3 ExpressPass 0.06 0.08 0.28 unstable 0.04 0.18 0.08 0.08 0.08 0.11 0.24 Homa 8.63 9.46 8.37 15.91 18.68 14.68 12.17 12.17 11.75 12.42 10.31 dcPIM 3.27 0.63 0.86 2.34 2.2 2.41 2.91 0.58 0.6 1.76 2.69 SIRD 0.76 0.81 0.75 2.26 1.67 1.39 0.79 0.81 0.83 1.12 1.51    Maximum ToR queuing across applied load levels (MB).
Table 5: Raw data used in Figure 5. Experiments in which the protocol is unable to deliver the specified throughput or network queuing grows infinitely (unstable) are excluded from the calculation of mean and range.
@@
Figure 13: Mean ToR queuing vs. achieved goodput. Configurations: Balanced (top), Core (middle), Incast (bottom).